OpenStudy (anonymous):

Find the Lim -> Inf. x^2+x+5 \ x^3 +1

OpenStudy (anonymous):

Since the limit is approaching inf the answer would be \[\frac{ 0+0+5 }{ 0+1 } = \] Correct?

OpenStudy (anonymous):

5. The answer i wrote was suppose to be 5

OpenStudy (mathmale):

GHHH: Careful here. If you're taking the limit of this rational expression as x approaches infinity, you have no justification for replacing x with zero. Want to give this further thought?

OpenStudy (mathmale):

Here is what I'd suggest: 1) Identify the highest power of x (it's x^3). 2) Divide EVERY term in numerator and denominator of this rational fraction by that highest power (x^3). 3) NOW, let x approach infinity. What's left?

OpenStudy (anonymous):

\[\frac{ \frac{ 1 }{ x }+ \frac{ 2 }{ x } + \frac{ 5 }{ x }}{ \frac{ x ^{3}{x ^{3}} }{ x ^{3}{x ^{3}}+ \frac{ 1 }{ x ^{3} }}}\] = 0?

OpenStudy (anonymous):

The second two X's on the bottom where suppose to be a fraction. I said zero because the top 3 turn into zeros becaus they're too large right? and the fractions are impossible to be found out?

OpenStudy (mathmale):

You have x^3 + 1 in the den. and are to divide both terms by x^3, producing 1 + 1/(x^3), right? I also suggested that you divide each term in the numerator by (x^3), which you have not yet done. Mind re-doing this division? Once you're done, again try taking the limit as x approaches infinity.

OpenStudy (mathmale):

Again, GHHH: Identify the HIGHEST POWER OF x (which is, in this case, x^3), and then divide EVERY TERM of the numerator and denom. by that highest power.

OpenStudy (anonymous):

\[\frac{ \frac{ 1 }{ x } + \frac{ x }{ x } + \frac{ 5 }{ x }}{ \frac{ x ^{3} }{ x ^{3} } +\frac{ 1 }{ x ^{3} } } = \frac{ 0 + 0 + 0 }{ 1 + 0 } = \frac{ 0 }{ 1 } = 0? Right? \]

OpenStudy (anonymous):

And I can't believe I missed the one in the dem, Sorry

OpenStudy (anonymous):

Sorry, not zero. Undefined! ***

OpenStudy (mathmale):

Please start over. ;) Divide every term, every single one of those terms, by x^3.

OpenStudy (mathmale):

Also, be certain that you're taking the limit as x approaches infinity (not zero).

OpenStudy (anonymous):

I am doing terrible. Sorry, okay so here it goes. \[\frac{ \frac{ x^2 }{ x^3 } +\frac{ x }{ x^3 }+\frac{ 5 }{ x^3 }}{ \frac{ x^3 }{ x^3 }+ \frac{ 1 }{ x^3 } } \] = \[\frac{ \frac{ 1 }{ x } +\frac{ x }{ x^2 }+\frac{ 5 }{ x^3 }}{ 1 + \frac{ 1 }{ x^3 } }\] = (Have an inf sign on the bottom = (extremely close to) zero. So by that logic, all number in on the top turn into zero because numbers out of inf. (Same applies to the 1/inf in the dom). So leaving us with 0/1. Which is undefined.

OpenStudy (anonymous):

Am I any closer?

OpenStudy (mathmale):

first of all, you are correct in stating that you're left with 0/1; however, 0/1 is zero, not undefined.

OpenStudy (anonymous):

Oh okay

OpenStudy (mathmale):

I'm going to take your original expression and demonstrate how I'd divide every term in it by x^3:\[\frac{ \frac{ x ^{2}+x+5 }{ x ^{3} } }{ \frac{ x ^{3}+1 }{ x ^{3} } }\]

OpenStudy (anonymous):

Okay

OpenStudy (mathmale):

That gives us: \[\frac{ \frac{ 1 }{ x }+\frac{ 1 }{ x ^{2} }+\frac{ 5 }{ x ^{3} } }{ 1 + \frac{ 1 }{ x ^{3}} }\]

OpenStudy (mathmale):

Now, Ghost, if we let x approach infinity, the entire numerator becomes 0 and the entire denominator becomes 1. Please be absolutely certain you understand what's happening here; if you don't, ask questions.

OpenStudy (mathmale):

so, if you end up with 0/1, that's = to 0, and 0 is the limit as x approaches infinity.

OpenStudy (anonymous):

Just for clarification: Since the denominator of the numerators are X (inf) they are going to equal Zero because any part of inf = zero because the part is too small. Hence, numerator = zero. The denominator = 1 because x^3 \ X^3 = 1. the other some had the same principle as above applied to it to make it zero. Leaving us with 0/1 = zero.

OpenStudy (anonymous):

If i do a 1 or 2 more, could you check it for me?

OpenStudy (mathmale):

I'd be glad to help you with another problem. Don't be offended, by my wanting to reword some of your "just for clarification" sentence for greater accuracy:

OpenStudy (anonymous):

no, of course not. Better wording is welcome.

OpenStudy (anonymous):

Clearer*

OpenStudy (mathmale):

"Every fraction in the numerator goes to zero as x goes to infinity because each of these fractions has a constant numerator, whereas the denominator of each fraction increases without bound." There's no "too small" here.

OpenStudy (mathmale):

All of the following go to zero as x increases to infinity:

OpenStudy (mathmale):

\[\frac{ 1 }{ x }, \frac{ 1 }{ x^2 }, \frac{ 1 }{ x^3 },....\] is the general principle that comes into play here.

OpenStudy (mathmale):

Try it: You can see for yourself that 1/x approaches 0 as x= 10, 100, 1000, etc.

OpenStudy (anonymous):

AH okay. Quick question then, Will there be a case where inf is the numerator with a constant as a denominator? As as for the try it, they increase by .0 each number. .1 -> .01 -> .001 -> etc.

OpenStudy (mathmale):

To answer your first question: sure. This fraction would have no limit. second question: actually, the fractions you've typed out here are DECREASING, which is my point exactly: as x goes to infinity, the fractions all get smaller, tending to 0 in the limit.